An unknown gaseous hydrocarbon consists of 85.63% carbon by mass. A 0.100-g sample of the gas occupies a volume of 0.0533 L at STP. What is the identity of the gas? A. C4H8 B. C3H6 C. C2H4 D. C5H10 E. CH2
Answer #1

% of carbon = 85.63%

% of Hydrogen = 100 - 85.63 = 14.37%

Calculate atomic ratio by dividing % composition by atomic mass.

Mass ratio of carbon = 85.63/12 = 7.13

Mass ratio of hydrogen = 14.37/1 = 14.37

Now calculate simplest ratio by dividing atomic ratio by smallest atomic ratio

Simplest ratio of carbon = 7.13/7.13 = 1

Simplest ratio of hydrogen = 14.37/7.13 = 2.015 = 2

Empirical formula of compound = CH2

According to avogadro's law 1 mole occupy volume 22.414 L

then 0.0533 L = 0.0533/22.414 = 0.002377978 mole

0.002377978 mole = 0.100 gm then 1 mole = 1 \times 0.100 / 0.002377978 =42.05 gm

molar mass of gas = 42.05 gm

Empirical formula mass = (12\times1) + (1\times2)  = 14

n = molecular formula mass / Empirical formula mass = 42.05/14 = 3

Molecular formula = n \times Empirical formula = 3 \times Empirical formula

Molecular formula of compound = C3H6

Answer - B. C3H6